OpenStudy (anonymous):

Solve the equation by completing the square. 4x^2-40x-12=0

OpenStudy (anonymous):

What's the first step to completing the square ?

OpenStudy (anonymous):

First divide the coefficient of x and square it..

OpenStudy (anonymous):

oh ok that makes sense

OpenStudy (anonymous):

but wait what are we dividing the coefficient of x by?

OpenStudy (anonymous):

We divide it by 2..

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

After squaring it add it to the equation..

OpenStudy (anonymous):

ok...

OpenStudy (anonymous):

To equal it to zero add a negative value also...

OpenStudy (anonymous):

i see

OpenStudy (mayankdevnani):

40x^2-40x-12=0 divide both sides by 4 to simplify. 10x^2-10x-3=0

OpenStudy (anonymous):

Yeah I forgot... @mayankdevnani has written a missing step.. :D

OpenStudy (mayankdevnani):

yaa

OpenStudy (anonymous):

oh simplifying it first... got it!

OpenStudy (anonymous):

Then the equation would be like this.. \[x^2-10x+25-28=0\]Solve this...

OpenStudy (anonymous):

 am i doing the right thing so far? ... sorry